OpenStudy (spicytaco1):

What are the solutions to the equation y² –1 = 24? Choose all answers that are correct. A. 5 B. 12.5 C. –5 D. –12.5 C?

OpenStudy (decentnabeel):

\[y^2-1=24\] \[\mathrm{Subtract\:}24\mathrm{\:from\:both\:sides}\] \[y^2-25=0\]

OpenStudy (decentnabeel):

@Spicytaco1 can you solved the Equation

OpenStudy (spicytaco1):

B?

OpenStudy (decentnabeel):

no

OpenStudy (decentnabeel):

Solve the Equation using factoring and completing square and quardaic formula... use any method

OpenStudy (spicytaco1):

a and c

OpenStudy (spicytaco1):

y = +5 and y = -5

OpenStudy (decentnabeel):

\[\mathrm{Solving\:with\:the\:Quadratic\:Formula}\] \[\mathrm{For\:a\:quadratic\:equation\:of\:the\:form\:}ax^2+bx+c=0\mathrm{\:the\:solutions\:are\:}\] \[\quad x_{1,\:2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[\mathrm{For\:}\quad a=1,\:b=0,\:c=-25:\quad y_{1,\:2}=\frac{-0\pm \sqrt{0^2-4\cdot \:1\left(-25\right)}}{2\cdot \:1}\] \[y=\frac{-0+\sqrt{0^2-4\cdot \:1\cdot \left(-25\right)}}{2\cdot \:1}=5\] \[y=\frac{-0-\sqrt{0^2-4\cdot \:1\cdot \left(-25\right)}}{2\cdot \:1}=-5\] \[y=5,\:y=-5\]

OpenStudy (spicytaco1):

can you help me with one more?

OpenStudy (decentnabeel):

yes

OpenStudy (spicytaco1):

ok

OpenStudy (spicytaco1):

What are the square roots of 484? A. −12 and 12 B. −18 and 18 C. −22 and 22 D. −242 and 242

OpenStudy (decentnabeel):

\[\sqrt{484}\] \[\mathrm{Factor\:the\:number:\:}\:484=22^2\] \[=\sqrt{22^2}\] \[\sqrt{22^2}=22\]

OpenStudy (spicytaco1):

Thanks :D

OpenStudy (decentnabeel):

Welcome :)